# Mixture of Experts Meets Prompt-Based Continual Learning

### v1

@@ Appendix A Statistical Suboptimality of Linear Gating Prefix MoE Model
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ are independent Gaussian noise variables such that $\mathbb{E}[{\varepsilon_{i}}|{\bm{X}}_{i}]=0$ and $\mathrm{Var}(\varepsilon_{i}|{\bm{X}}_{i})=\nu^{2}$ for all $1\leq i\leq n$. Additionally, we assume that ${\bm{X}}_{1},{\bm{X}}_{2},\ldots,{\bm{X}}_{n}$ are i.i.d. samples from some probability distribution $\mu$. Motivated by linear gating prefix MoE model (11), the regression function $f_{G_{*}}(\cdot)$ in equation (20) admits the form of the linear gating prefix MoE model with pretrained $N$ experts and $L$ unknown experts, namely
@@ Appendix B Proof of Theoretical Results
$\mathcal{F}_{k}(\Theta,\delta):=\left\{g\in\mathcal{F}_{k}(\Theta):\|g-g_{G_{*}}\|_{L^{2}(\mu)}\leq\delta\right\}.$
@@
$H_{B}(\varepsilon,\mathcal{F}_{L^{\prime}}(\Theta),\|.\|_{L^{2}(\mu)})\lesssim\log(1/\varepsilon),$ (31)
for any $20<\varepsilon\leq 1/$ 2. Indeed, for any function $g_{G}\in\mathcal{F}_{L^{\prime}}(\Theta)$, since the expert functions are bounded, we obtain that $h({\bm{X}},\eta)\leq M$ for all ${\bm{X}}$ where $M$ is a bounded constant of the expert functions. Let $\tau\leq\varepsilon$ and $\{\pi_{1},\ldots,\pi_{\bar{N}}\}$ be the $\tau$-cover under the $L^{2}$ norm of the set $\mathcal{F}_{L^{\prime}}(\Theta)$ where $\bar{N}:={N}(\tau,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L^{2}(\mu)})$ is the $\eta$-covering number of the metric space $(\mathcal{F}_{k}(\Theta),\|\cdot\|_{L^{2}(\mu)})$. Then, we construct the brackets of the form $[L_{i}({\bm{X}}),U_{i}({\bm{X}})]$ for all $i\in[\bar{N}]$ as follows:
$L_{i}(x)$ $:=\max\{\pi_{i}({\bm{X}})-\tau,0\},$
$U_{i}(x)$ $:=\max\{\pi_{i}({\bm{X}})+\tau,M\}.$
From the above construction, we can validate that $\mathcal{F}_{L^{\prime}}(\Theta)\subset\cup_{i=1}^{\bar{N}}[L_{i}({\bm{X}}),U_{i}({\bm{X}})]$ and $U_{i}({\bm{X}})-L_{i}({\bm{X}})\leq 2\min\{2\tau,M\}$. Therefore, it follows that
$\|U_{i}-L_{i}\|_{L_{2}(\mu)}^{2}=\int(U_{i}-L_{i})^{2}\mathrm{d}\mu({\bm{X}})\leq\int 16\tau^{2}\mathrm{d}\mu({\bm{X}})=16\tau^{2},$
which implies that $\|U_{i}-L_{i}\|_{L_{2}(\mu)}\leq 4\tau$. By definition of the bracketing entropy, we deduce that
$H_{B}(4\tau,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L_{2}(\mu)})\leq\log N=\log{N}(\tau,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L^{2}(\mu)}).$ (32)
Therefore, we need to provide an upper bound for the covering number $\bar{N}$. In particular, we denote $\Delta:=\{(\beta_{1},\beta_{0})\in\mathbb{R}^{Nd\times Nd}\times\mathbb{R}^{Nd}\times\mathbb{R}:(\beta_{1},\beta_{0},\eta)\in\Theta\}$ and $\Omega:=\{\eta\in\mathbb{R}^{q}:(\beta_{1},\beta_{0},\eta)\in\Theta\}$. Since $\Theta$ is a compact set, $\Delta$ and $\Omega$ are also compact. Therefore, we can find $\tau$-covers $\Delta_{\tau}$ and ${\Omega}_{\tau}$ for $\Delta$ and $\Omega$, respectively. We can check that
$|\Delta_{\tau}|\leq\mathcal{O}_{P}(\tau^{-(Nd+1)L^{\prime}}),\quad|\Omega_{\tau}|\lesssim\mathcal{O}_{P}(\tau^{-qL^{\prime}}).$
@@
Here, $\overline{\eta}_{i}\in{\Omega}_{\tau}$ such that $\overline{\eta}_{i}$ is the closest to $\eta_{i}$ in that set, while $(\overline{\beta}_{1i},\overline{\beta}_{0i})\in\Delta_{\tau}$ is the closest to $(\beta_{1i},\beta_{0i})$ in that set. From the above formulations, we get that
$\|g_{G}-g_{\check{G}}\|_{L_{2}(\mu)}^{2}$
$=\int\Bigg{[}\sum_{j=1}^{L^{\prime}}\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})}{\sum_{i^{\prime}=1}^{N}\exp({\bm{X}}^{\top}B^{0}_{i^{\prime}}{\bm{X}}+c^{0}_{i^{\prime}})+\sum_{j^{\prime}=1}^{L^{\prime}}\exp(\beta_{1j^{\prime}}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j^{\prime}}^{\top}x)+\beta_{0j^{\prime}})}$
×$\hskip 227.62204pt\times\big{(}h({\bm{X}},\eta_{j})-h({\bm{X}},\overline{\eta}_{j})\big{)}\Bigg{]}^{2}~{}\mathrm{d}\mu({\bm{X}})$
$\leq L^{\prime}\int\sum_{j=1}^{L^{\prime}}\Bigg{[}\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})}{\sum_{i^{\prime}=1}^{N}\exp({\bm{X}}^{\top}B^{0}_{i^{\prime}}{\bm{X}}+c^{0}_{i^{\prime}})+\sum_{j^{\prime}=1}^{L^{\prime}}\exp(\beta_{1j^{\prime}}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j^{\prime}}^{\top}{\bm{X}})+\beta_{0j^{\prime}})}$
×$\hskip 227.62204pt\times\big{(}h({\bm{X}},\eta_{j})-h({\bm{X}},\overline{\eta}_{j})\big{)}\Bigg{]}^{2}~{}\mathrm{d}\mu({\bm{X}})$
$\leq L^{\prime}\int\sum_{j=1}^{L^{\prime}}~{}[h({\bm{X}},\eta_{j})-h({\bm{X}},\overline{\eta}_{j})]^{2}~{}\mathrm{d}\mu({\bm{X}})$
$\leq L^{\prime}\int\sum_{j=1}^{L^{\prime}}~{}[L_{1}\cdot\|\eta_{j}-\overline{\eta}_{j}\|]^{2}~{}\mathrm{d}\mu({\bm{X}})$
$\leq(L^{\prime}L_{1}\tau)^{2},$
which indicates that $\|g_{G}-g_{\check{G}}\|_{L_{2}(\mu)}\lesssim\tau$. Here, the second inequality is according to the Cauchy-Schwarz inequality, the third inequality occurs as the softmax weight is bounded by 1, and the fourth inequality follows from the fact that the expert $h(x,\cdot)$ is a Lipschitz function with Lipschitz constant $L_{1}$. Next, let us denote
@@
$\|g_{\check{G}}-g_{\overline{G}}\|_{L_{2}(\mu)}^{2}=\int\Big{[}\frac{1}{D}\Big{(}\sum_{i=1}^{N}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})+\sum_{j=1}^{L^{\prime}}\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})h({\bm{X}},\overline{\eta}_{j})\Big{)}$
$+\frac{1}{\overline{D}}\Big{(}\sum_{i=1}^{N}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})+\sum_{j=1}^{L^{\prime}}\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\overline{\beta}_{1j}^{\top}{\bm{X}})+\overline{\beta}_{0j})h({\bm{X}},\overline{\eta}_{j})\Big{)}\Big{]}^{2}\mathrm{d}\mu({\bm{X}})$
$\leq\frac{1}{2}\int\Bigg{\{}\Big{[}\sum_{i=1}^{N}\Big{(}\frac{\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})}{D}-\frac{\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})}{\overline{D}}\Big{)}h({\bm{X}},\eta_{i}^{0})\Big{]}^{2}$
$\hskip 56.9055pt+\Big{[}\sum_{j=1}^{L^{\prime}}\Big{(}\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})}{D}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})}{\overline{D}}\Big{)}h({\bm{X}},\overline{\eta}_{j})\Big{]}^{2}\Bigg{\}}\mathrm{d}\mu({\bm{X}})$
$\leq\frac{N}{2}\Big{(}\frac{1}{D}-\frac{1}{\overline{D}}\Big{)}^{2}\int\sum_{i=1}^{N}\Big{[}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})\Big{]}^{2}\mathrm{d}\mu({\bm{X}})$
$\hskip 42.67912pt+\frac{L^{\prime}}{2}\int\sum_{j=1}^{L^{\prime}}\Bigg{[}\Big{(}\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})}{D}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})}{\overline{D}}\Big{)}h({\bm{X}},\overline{\eta}_{j})\Bigg{]}^{2}~{}\mathrm{d}\mu({\bm{X}}).$ (33)
@@
$=\Big{|}\sum_{j^{\prime}=1}^{L^{\prime}}\Big{[}\exp(\beta_{1j^{\prime}}^{\top}{\bm{X}}+\sigma(\beta_{1j^{\prime}}^{\top}{\bm{X}})+\beta_{0j^{\prime}})-\exp(\overline{\beta}_{1j^{\prime}}^{\top}{\bm{X}}+\sigma(\overline{\beta}_{1j^{\prime}}^{\top}{\bm{X}})+\overline{\beta}_{0j^{\prime}})\Big{]}\Big{|}$
$\lesssim\sum_{j=1}^{L^{\prime}}\Big{[}\|\beta_{1j}-\overline{\beta}_{1j}\|\cdot\|x\|+|\beta_{0j}-\overline{\beta}_{0j}|\Big{]}$
$\leq k\tau(B+1).$
@@
$\frac{k_{0}}{2}\Big{(}\frac{1}{D}-\frac{1}{\overline{D}}\Big{)}^{2}\int\sum_{i=1}^{N}\Big{[}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})\Big{]}^{2}\mathrm{d}\mu({\bm{X}})\lesssim\frac{1}{2}N[L^{\prime}\tau(B+1)]^{2}.$ (34)
@@
$\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})}{D}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})}{\overline{D}}$
$=$ $~{}\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})\Big{(}\frac{1}{D}-\frac{1}{\overline{D}}\Big{)}+\frac{1}{\overline{D}}\Big{[}\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})-\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})\Big{]}.$
Since
$\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})\Big{(}\frac{1}{D}-\frac{1}{\overline{D}}\Big{)}$ $\lesssim\frac{1}{D}-\frac{1}{\overline{D}}\lesssim L^{\prime}\tau(B+1),$
$\frac{1}{\overline{D}}\Big{[}\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})-\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})\Big{]}$ $\lesssim\Big{[}\|\beta_{1j}-\overline{\beta}_{1j}\|\cdot\|{\bm{X}}\|+|\beta_{0j}-\overline{\beta}_{0j}|\Big{]}\leq\tau(B+1),$
it follows that
$\frac{L^{\prime}}{2}\int\sum_{j=1}^{L^{\prime}}\Bigg{[}\Big{(}\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\beta_{0j})}{D}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\overline{\beta}_{0j})}{\overline{D}}\Big{)}h(x,\overline{\eta}_{j})\Bigg{]}^{2}~{}\mathrm{d}\mu(x)\lesssim\frac{1}{2}(L^{\prime})^{2}M^{2}[\tau(B+1)]^{2}$ (35)
From equations (33), (34) and (35), we obtain that
$\|g_{\check{G}}-g_{\overline{G}}\|_{L_{2}(\mu)}\lesssim\tau.$
According to the triangle inequality, we have
$\|g_{G}-g_{\overline{G}}\|_{L_{2}(\mu)}\leq\|g_{G}-g_{\check{G}}\|_{L_{2}(\mu)}+\|g_{\check{G}}-g_{\overline{G}}\|_{L_{2}(\mu)}\lesssim\tau.$
@@
${N}(\tau,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L_{2}(\mu)})\leq|\Delta_{\tau}|\times|\Omega_{\tau}|\leq\mathcal{O}(n^{-(Nd+1)L^{\prime}})\times\mathcal{O}(n^{-qL^{\prime}})\leq\mathcal{O}(n^{-(Nd+1+q)L^{\prime}}).$ (36)
@@
$H_{B}(4\tau,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L_{2}(\mu)})\lesssim\log(1/\tau).$
Let $\tau=\varepsilon/4$, then we obtain that
$H_{B}(\varepsilon,\mathcal{F}_{L^{\prime}}(\Theta),\|.\|_{L^{2}(\mu)})\lesssim\log(1/\varepsilon).$
@@
$u$$\in\left\{{\bm{X}}^{\top}B^{0}_{i^{\prime}}{\bm{X}}+c^{0}_{i^{\prime}},\ (\beta_{1j^{\prime}})^{\top}{\bm{X}}+\sigma((\beta_{1j^{\prime}})^{\top}{\bm{X}})+\beta_{0j^{\prime}}:i^{\prime}\in[N],j^{\prime}\in[L^{\prime}]\right\},$
@@
Since $g_{G}({\bm{X}})=g_{G_{*}}({\bm{X}})$ for almost every $x$, we have

### v2

@@ Appendix A Statistical Suboptimality of Linear Gating Prefix MoE Model
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ are independent Gaussian noise variables such that $\mathbb{E}[{\varepsilon_{i}}|{\bm{X}}_{i}]=0$ and $\mathrm{Var}(\varepsilon_{i}|{\bm{X}}_{i})=\nu^{2}$ for all $1\leq i\leq n$. Additionally, we assume that ${\bm{X}}_{1},{\bm{X}}_{2},\ldots,{\bm{X}}_{n}$ are i.i.d. samples from some probability distribution $\mu$. Motivated by linear gating prefix MoE model (11), the regression function $f_{G_{*}}(\cdot)$ in equation (23) admits the form of the linear gating prefix MoE model with pre-trained $N$ experts and $L$ unknown experts, namely
@@ Appendix B Proof of Theoretical Results
$\mathcal{F}_{L^{\prime}}(\Theta,\delta):=\left\{g\in\mathcal{F}_{L^{\prime}}(\Theta):\|g-g_{G_{*}}\|_{L^{2}(\mu)}\leq\delta\right\}.$
@@
$H_{B}(\varepsilon,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L^{2}(\mu)})\lesssim\log(1/\varepsilon),$ (34)
for any $20<\varepsilon\leq 1/$ 2. Indeed, for any function $g_{G}\in\mathcal{F}_{L^{\prime}}(\Theta)$, since the expert functions are bounded, we obtain that $h({\bm{X}},\eta)\leq M$ for all ${\bm{X}}$ where $M$ is a bounded constant of the expert functions. Let $\tau\leq\varepsilon$ and $\{\pi_{1},\ldots,\pi_{\bar{N}}\}$ be the $\zeta$-cover under the $L^{\infty}$ norm of the set $\mathcal{F}_{L^{\prime}}(\Theta)$ where $\bar{N}:={N}(\zeta,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L^{\infty}})$ is the $\eta$-covering number of the metric space $(\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L^{\infty}})$. Then, we construct the brackets of the form $[L_{i}({\bm{X}}),U_{i}({\bm{X}})]$ for all $i\in[\bar{N}]$ as follows:
$L_{i}(x)$ $:=\max\{\pi_{i}({\bm{X}})-\zeta,0\},$
$U_{i}(x)$ $:=\max\{\pi_{i}({\bm{X}})+\zeta,M\}.$
From the above construction, we can validate that $\mathcal{F}_{L^{\prime}}(\Theta)\subset\cup_{i=1}^{\bar{N}}[L_{i}({\bm{X}}),U_{i}({\bm{X}})]$ and $U_{i}({\bm{X}})-L_{i}({\bm{X}})\leq\min\{2\zeta,M\}$. Therefore, it follows that
$\|U_{i}-L_{i}\|_{L_{2}(\mu)}^{2}=\int(U_{i}-L_{i})^{2}\mathrm{d}\mu({\bm{X}})\leq\int 4\zeta^{2}\mathrm{d}\mu({\bm{X}})=4\zeta^{2},$
which implies that $\|U_{i}-L_{i}\|_{L_{2}(\mu)}\leq 2\zeta$. By definition of the bracketing entropy, we deduce that
$H_{B}(2\zeta,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L_{2}(\mu)})\leq\log N=\log{N}(\zeta,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L^{\infty}}).$ (35)
Therefore, we need to provide an upper bound for the covering number $\bar{N}$. In particular, we denote $\Delta:=\{(\beta_{1},\beta_{0})\in\mathbb{R}^{Nd\times Nd}\times\mathbb{R}^{Nd}\times\mathbb{R}:(\beta_{1},\beta_{0},\eta)\in\Theta\}$ and $\Omega:=\{\eta\in\mathbb{R}^{q}:(\beta_{1},\beta_{0},\eta)\in\Theta\}$. Since $\Theta$ is a compact set, $\Delta$ and $\Omega$ are also compact. Therefore, we can find $\zeta$-covers $\Delta_{\zeta}$ and ${\Omega}_{\zeta}$ for $\Delta$ and $\Omega$, respectively. We can check that
$|\Delta_{\zeta}|\leq\mathcal{O}_{P}(\tau^{-(Nd+1)L^{\prime}}),\quad|\Omega_{\zeta}|\lesssim\mathcal{O}_{P}(\tau^{-qL^{\prime}}).$
@@
Here, $\overline{\eta}_{i}\in{\Omega}_{\zeta}$ such that $\overline{\eta}_{i}$ is the closest to $\eta_{i}$ in that set, while $(\overline{\beta}_{1i},\overline{\beta}_{0i})\in\Delta_{\zeta}$ is the closest to $(\beta_{1i},\beta_{0i})$ in that set. From the above formulations, we get that
$\|g_{G}-g_{\check{G}}\|_{L^{\infty}}$
$=\sup_{{\bm{X}}\in\mathcal{X}}~{}\sum_{j=1}^{L^{\prime}}\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})\cdot|h({\bm{X}},\eta_{j})-h({\bm{X}},\overline{\eta}_{j})|}{\sum_{i^{\prime}=1}^{N}\exp({\bm{X}}^{\top}B^{0}_{i^{\prime}}{\bm{X}}+c^{0}_{i^{\prime}})+\sum_{j^{\prime}=1}^{L^{\prime}}\exp(\beta_{1j^{\prime}}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j^{\prime}}^{\top}x)+\beta_{0j^{\prime}})}$
$\leq\sum_{j=1}^{L^{\prime}}\sup_{{\bm{X}}\in\mathcal{X}}~{}\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})\cdot|h({\bm{X}},\eta_{j})-h({\bm{X}},\overline{\eta}_{j})|}{\sum_{i^{\prime}=1}^{N}\exp({\bm{X}}^{\top}B^{0}_{i^{\prime}}{\bm{X}}+c^{0}_{i^{\prime}})+\sum_{j^{\prime}=1}^{L^{\prime}}\exp(\beta_{1j^{\prime}}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j^{\prime}}^{\top}{\bm{X}})+\beta_{0j^{\prime}})}$
$\leq\sum_{j=1}^{L^{\prime}}\sup_{{\bm{X}}\in\mathcal{X}}~{}|h({\bm{X}},\eta_{j})-h({\bm{X}},\overline{\eta}_{j})|$
$\leq\sum_{j=1}^{L^{\prime}}\sup_{{\bm{X}}\in\mathcal{X}}~{}[L_{1}({\bm{X}})\cdot\|\eta_{j}-\overline{\eta}_{j}\|]$
$\lesssim L^{\prime}\zeta\lesssim\zeta.$
Here, the second inequality occurs as the softmax weight is bounded by one, and the third inequality follows from the fact that the expert $h({\bm{X}},\cdot)$ is a Lipschitz function with some Lipschitz constant $L_{1}({\bm{X}})>0$. Next, let us denote
@@
$\|g_{\check{G}}-g_{\overline{G}}\|_{L^{\infty}}$
$=\sup_{{\bm{X}}\in\mathcal{X}}\Bigg{|}\frac{1}{D}\Big{(}\sum_{i=1}^{N}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})+\sum_{j=1}^{L^{\prime}}\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})h({\bm{X}},\overline{\eta}_{j})\Big{)}$
$\quad-\frac{1}{\overline{D}}\Big{(}\sum_{i=1}^{N}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})+\sum_{j=1}^{L^{\prime}}\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\overline{\beta}_{1j}^{\top}{\bm{X}})+\overline{\beta}_{0j})h({\bm{X}},\overline{\eta}_{j})\Big{)}\Bigg{|}$
$\leq\Big{|}\frac{1}{D}-\frac{1}{\overline{D}}\Big{|}\cdot\sum_{i=1}^{N}\sup_{{\bm{X}}\in\mathcal{X}}~{}\Big{|}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})\Big{|}$
$\quad+\sum_{j=1}^{L^{\prime}}\sup_{{\bm{X}}\in\mathcal{X}}~{}\Bigg{|}\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})}{D}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\overline{\beta}_{1j}^{\top}{\bm{X}})+\overline{\beta}_{0j})}{\overline{D}}\Bigg{|}\cdot|h({\bm{X}},\overline{\eta}_{j})|.$ (36)
@@
$\leq\sum_{j^{\prime}=1}^{L^{\prime}}\Big{|}\exp(\beta_{1j^{\prime}}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j^{\prime}}^{\top}{\bm{X}})+\beta_{0j^{\prime}})-\exp(\overline{\beta}_{1j^{\prime}}^{\top}{\bm{X}}+\alpha\sigma(\tau\overline{\beta}_{1j^{\prime}}^{\top}{\bm{X}})+\overline{\beta}_{0j^{\prime}})\Big{|}$
$\lesssim\sum_{j^{\prime}=1}^{L^{\prime}}\Big{|}(\beta_{1j}-\overline{\beta}_{1j})^{\top}{\bm{X}}+\alpha[\sigma(\tau\beta_{1j^{\prime}}^{\top}{\bm{X}})-\sigma(\tau\overline{\beta}_{1j^{\prime}}^{\top}{\bm{X}})]+(\beta_{0j}-\overline{\beta}_{0j})\Big{|}$
$\leq\sum_{j^{\prime}=1}^{L^{\prime}}|(\beta_{1j}-\overline{\beta}_{1j})^{\top}{\bm{X}}|+|\alpha|\cdot|\sigma(\tau\beta_{1j^{\prime}}^{\top}{\bm{X}})-\sigma(\tau\overline{\beta}_{1j^{\prime}}^{\top}{\bm{X}})|+|\beta_{0j}-\overline{\beta}_{0j}|$
$\lesssim\sum_{j=1}^{L^{\prime}}\Big{[}\|\beta_{1j}-\overline{\beta}_{1j}\|\cdot\|{\bm{X}}\|+|\alpha\tau|\cdot\|\beta_{1j}-\overline{\beta}_{1j}\|\cdot\|{\bm{X}}\|+|\beta_{0j}-\overline{\beta}_{0j}|\Big{]}$
$\leq L^{\prime}(B+|\alpha\tau|B+1)\zeta\lesssim\zeta.$
@@
$\Big{|}\frac{1}{D}-\frac{1}{\overline{D}}\Big{|}\cdot\sum_{i=1}^{N}\sup_{{\bm{X}}\in\mathcal{X}}~{}\Big{|}\exp({\bm{X}}^{\top}B_{i}^{0}{\bm{X}}+c_{i}^{0})h({\bm{X}},\eta_{i}^{0})\Big{|}\lesssim\zeta.$ (37)
@@
$\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})}{D}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\overline{\beta}_{1j}^{\top}{\bm{X}})+\overline{\beta}_{0j})}{\overline{D}}$
$=$ $~{}\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})\Big{(}\frac{1}{D}-\frac{1}{\overline{D}}\Big{)}$
$\qquad+\frac{1}{\overline{D}}\Big{[}\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})-\exp(\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\overline{\beta}_{1j}^{\top}{\bm{X}})+\overline{\beta}_{0j}))\Big{]}.$
Since both the input space and the parameter space are bounded, we have
$\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})\Big{(}\frac{1}{D}-\frac{1}{\overline{D}}\Big{)}\lesssim\frac{1}{D}-\frac{1}{\overline{D}}\lesssim\zeta,$
$\frac{1}{\overline{D}}\Big{[}\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})-\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\overline{\beta}_{1j}^{\top}{\bm{X}})+\overline{\beta}_{0j})$
$\hskip 199.16928pt\lesssim(B+|\alpha\tau|B+1)\zeta\lesssim\zeta,$
which yields that
$\sum_{j=1}^{L^{\prime}}\sup_{{\bm{X}}\in\mathcal{X}}~{}\Bigg{|}\frac{\exp(\beta_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\beta_{1j}^{\top}{\bm{X}})+\beta_{0j})}{D}-\frac{\exp(\overline{\beta}_{1j}^{\top}{\bm{X}}+\alpha\sigma(\tau\overline{\beta}_{1j}^{\top}{\bm{X}})+\overline{\beta}_{0j})}{\overline{D}}\Bigg{|}\cdot|h({\bm{X}},\overline{\eta}_{j})|$
$\hskip 227.62204pt\lesssim\zeta~{}\sum_{j=1}^{L^{\prime}}\sup_{{\bm{X}}\in\mathcal{X}}~{}|h({\bm{X}},\overline{\eta}_{j})|\lesssim\zeta.$ (38)
From equations (36), (37) and (B.1), we obtain that $\|g_{\check{G}}-g_{\overline{G}}\|_{L^{\infty}}\lesssim\zeta$. According to the triangle inequality, we have
$\|g_{G}-g_{\overline{G}}\|_{L^{\infty}}\leq\|g_{G}-g_{\check{G}}\|_{L^{\infty}}+\|g_{\check{G}}-g_{\overline{G}}\|_{L^{\infty}}\lesssim\zeta.$
@@
${N}(\zeta,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L^{\infty}})\leq|\Delta_{\zeta}|\times|\Omega_{\zeta}|\leq\mathcal{O}(n^{-(Nd+1)L^{\prime}})\times\mathcal{O}(n^{-qL^{\prime}})\leq\mathcal{O}(n^{-(Nd+1+q)L^{\prime}}).$ (39)
@@
$H_{B}(2\zeta,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L_{2}(\mu)})\lesssim\log(1/\tau).$
Let $\zeta=\varepsilon/2$, then we obtain that
$H_{B}(\varepsilon,\mathcal{F}_{L^{\prime}}(\Theta),\|\cdot\|_{L^{2}(\mu)})\lesssim\log(1/\varepsilon).$
@@
$u$$\in\left\{{\bm{X}}^{\top}B^{0}_{i^{\prime}}{\bm{X}}+c^{0}_{i^{\prime}},\ (\beta_{1j^{\prime}})^{\top}{\bm{X}}+\alpha\sigma(\tau(\beta_{1j^{\prime}})^{\top}{\bm{X}})+\beta_{0j^{\prime}}:i^{\prime}\in[N],j^{\prime}\in[L^{\prime}]\right\},$
@@
Since $g_{G}({\bm{X}})=g_{G_{*}}({\bm{X}})$ for almost every ${\bm{X}}$, we have